# Dimensional crossover of class D real-space topological invariants

### v1

@@ III Spectral localizer
The 1D local topological invariant in class D is defined by first transforming $\mathcal{H}$ to be in a basis in which it is purely imaginary, so that $\mathcal{H}^{\mathsf{T}}=-\mathcal{H}$. Such a basis must exist due to the system’s particle-hole symmetry. In this basis for $\mathcal{H}$, and noting the choice of Pauli matrices used in Eq. (7), the 1D spectral localizer is both block off-diagonal and real symmetric, $(L^{(1D)}_{\bm{\lambda}})^{\dagger}=(L^{(1D)}_{\bm{\lambda}})^{\mathsf{T}}=L^{(1D)}_{\bm{\lambda}}$. Thus, all of its eigenvalues come in pairs, and when $L^{(1D)}_{\bm{\lambda}}$ is invertible, its matrix homotopy can be classified using only one of its off-diagonal blocks (which removes this spectral pairing) as
$z_{2,\bm{\lambda}}=\text{sign}\left[\text{det}\left\{\kappa(X-x\mathcal{I})-i(H-E\mathcal{I})\right\}\right],$ (8)
see Theorem D1 in Ref. [62] for an accessible explanation.
In the next section, we calculate the 1D and 2D topological invariants for a chain and an island, respectively, to define a topological phase diagram. We then characterize the topological transitions induced by the change in dimension, as a Shiba island is deformed into a Shiba chain.
Figure 5: (a)-(e): Arrangements of magnetic adatoms interpolating from the two-dimensional circular island to the one-dimensional chain geometries. (f)-(j): 1D (black solid line) and 2D (dashed orange lines) localizer spectrum $\Sigma(L_{\bm{\lambda}})$ as a function of position $x$ for $y=N/2$. (k)-(o): Local Chern number, $|C_{\bm{\lambda}}^{\textrm{L}}|$, (orange squares) and 2D localizer gap $\text{min}\{\left|\Sigma(L^{(2D)}_{\bm{\lambda}})\right|\}$ (solid gray line) as a function of position $x$ for $y=N/2$. (p)-(t): Local $z_{2,\bm{\lambda}}$ invariant (black circles) and 1D localizer gap $\text{min}\{\left|\Sigma(L^{(1D)}_{\bm{\lambda}})\right|\}$ (solid gray line) as a function of position $x$ for $y=N/2$. The parameters are $N=20$, $\mu=1.0$, $\alpha=0.45$, $\Delta=1.2$, and $JS=2.579$. The spectral localizer calculations use $\kappa=0.05(t/l)$.
IV Results
@@
Leveraging the local topological invariants for Class D systems, we explore the Shiba lattice’s parameter space for 1D and 2D geometries and consider values previously used in the literature for this model [63, 55, 64, 56]. To this end, we consider two geometries, a circular island, and a chain, and compute the corresponding local topological invariants (the Chern number $C_{\bm{\lambda}}^{\textrm{L}}$ and $z_{2,\bm{\lambda}}$, respectively) at the system’s center. The results are plotted in Fig. 3. Using the same parameters for both geometries, we find regions where the Shiba island and chain are both topological, regions where the Shiba island is topological and the chain is trivial, regions where the island is trivial and the chain is topological, and regions where both geometries are trivial. Note that topological transitions not captured in the diagram can occur as the geometry changes from 2D to 1D. For example, the transition from $|C_{\bm{\lambda}}^{\textrm{L}}|=2$ to $z_{2,\bm{\lambda}}=-1$ as the circular island is flattened occurs via a change to $|C_{\bm{\lambda}}^{\textrm{L}}|=1$ for a narrow island, as we discuss later.
Next, we consider representative cases in the parameter space and study the topological invariants as the Shiba geometry transitions from two-dimensional to one-dimensional.
IV.1 Topological invariant dimensional crossover
@@
Figure 6: Length of the cut along $x$ for $y=N/2$, as shown in Fig. (4), for which the local topological invariant is nontrivial, and the local gap is larger than $0.02t$, $d^{top}$, divided by the number of sites along the cut, $d^{tot}$. The parameters for panel (a) are the same as in Fig. 4. In panel (b), we show the case Fig. 9 with parameters $N=20$, $\mu=0$, $\alpha=0.8$, $\Delta=1.2$, and $JS=2$. Panel (c) shows the case in Fig. 5, the parameters are $N=20$, $\mu=1.0$, $\alpha=0.45$, $\Delta=1.2$, and $JS=2.579$. The spectral localizer calculations use $\kappa=0.05(t/l)$.
First, we consider the case where both the Shiba island and chain are topological with $|C_{\bm{\lambda}}^{\textrm{L}}|=1$ and $z_{2,\bm{\lambda}}=-1$, respectively. In Fig. 4, top row, we plot the set of magnetic adatom geometries considered, which interpolate between a Shiba circular island [(a)] and a Shiba chain [(e)] by reducing the number of magnetic adatoms along the y-direction. In the second row, we plot the spectra of the 1D (black solid lines) and 2D (dashed orange lines) spectral localizers, $\Sigma(L_{\bm{\lambda}}(\bm{X},\mathcal{H}))$, computed as a function of position $x$, for $y=N/2$ and zero energy.
@@
For the Shiba island, panel (f), $\Sigma(L^{(2D)}_{\bm{\lambda}}(\bm{X},\mathcal{H}))$ crosses zero at the edge of the magnetic adatom island, transitioning from trivial $C_{\bm{\lambda}}^{\textrm{L}}=0$ to topological with $|C_{\bm{\lambda}}^{\textrm{L}}|=1$ (panel (k)). The finite localizer gap is maximal at the center of the island. On the other hand, the spectrum of the 1D spectral localizer, $\Sigma(L^{(1D)}_{\bm{\lambda}}(\bm{X},\mathcal{H}))$, displays a much smaller local gap at the location of magnetic adatoms. While there are sites where the topological local marker is topological, $z_{2,\bm{\lambda}}=-1$, there is no local gap protection against variations in the Hamiltonian $\delta\mathcal{H}$ or position operators $\delta X$. Therefore, the 1D topology is trivial, and the system is characterized by a two-dimensional topology via the local Chern number, $C_{\bm{\lambda}}^{\textrm{L}}$.
@@
We now reduce the number of magnetic adatoms along the y-direction, flattening the island, as shown in panel (b). Similar trends in the local markers and the localizer gaps are obtained. However, the number of sites with a nontrivial Chern number is reduced compared with the circular island’s case. Additionally, the 2D localizer gap is now smaller. The trend continues for a flatter island, as shown in panel (c). However, for the parameters considered, the three-atom-tall island exhibits a behavior change. The Chern number is now nonzero only at locations where the 2D localizer gap vanishes, panel (n), while there are sites with a nontrivial $z_{2,\bm{\lambda}}$ invariant protected by a finite 1D localizer gap, panel (s). This marks the transition from 2D to 1D topology in the system. For the Shiba chain, panel (e), the $z_{2,\bm{\lambda}}$ invariant changes from trivial to nontrivial at the edge of the chain. As the Shiba island reduces to a chain, $\Sigma(L^{(2D)}_{\bm{\lambda}=(x,y=N/2,E=0)}(\bm{X},\mathcal{H}))$ approaches $\Sigma(L^{(1D)}_{\bm{\lambda}=(x,E=0)}(\bm{X},\mathcal{H}))$.
@@
We now consider a set of parameters that leads to a topological island with $|C_{\bm{\lambda}}^{\textrm{L}}|=2$ transitioning into a topological chain. We show our results in Fig. 5. There are two main differences with respect to the previous example we discussed. First, there is a topological transition from $|C_{\bm{\lambda}}^{\textrm{L}}|=2$ to $|C_{\bm{\lambda}}^{\textrm{L}}|=1$ when the Shiba island shrinks from seven to five atoms tall. Second, for the three-atom-tall island, there is a coexistence of 2D and 1D topology with $|C_{\bm{\lambda}}^{\textrm{L}}|=1$ and $z_{2,\bm{\lambda}}=-1$, respectively. However, as we will discuss later, when including disorder in the chemical potential, the two-dimensional local gap is more sensitive to disorder than the 1D local gap.
@@
Finally, in Appendix A, we show plots for a topological island with $|C_{\bm{\lambda}}^{\textrm{L}}|=2$ transitioning into the trivial chain. In this case, the number of sites with nontrivial topology decreases together with the local gap protection, until it becomes trivial in both 1D and 2D. We present a graphical summary of our calculations in Fig. 6, where we plot the number of sites with a topological local marker, $d^{top}$, divided by the total number of sites on the $x$-axis, $d^{tot}$, both for 1D and 2D, as a function of the island’s width to length ratio. We only consider sites topological if the local gap is larger than $0.02t$. The transition from 2D topology to 1D topology, as characterized by the spectral localizer, is given by the crossing of the traces.
@@
Figure 7: (a) 2D localizer gap for a disordered circular Shiba island (as shown in the inset) along position $x$ and $y=N/2$, with disorder strength $W/t=1.2$. The dots show the average over 100 disorder realizations, while the shaded area corresponds to ±$\pm 1$ standard deviations. Panel (b) shows the corresponding local maker, $|C_{\bm{\lambda}}^{\textrm{L}}|$, which is trivial. (c) 1D localizer gap for a disordered Shiba chain (as shown in the inset) along position $x$ and $y=N/2$, also for $W/t=1.2$, corresponding to the chain’s position. Panel (d) shows the corresponding local marker, $z_{2,\bm{\lambda}}$, which remains topological. The spectral localizer calculations use $\kappa=0.05(t/l)$.
IV.2 Effects of disorder
@@
To test the robustness of our result, we consider disorder in the local chemical potential. The additional term we include in the Hamiltonian is given by
$\mathcal{H}_{dis}=-\sum_{\mathbf{r},\sigma}\delta\mu(\mathbf{r})c^{\dagger}_{\mathbf{r}\sigma}c_{\mathbf{r}\sigma},$ (9)
@@
where $\delta\mu(\mathbf{r})$ is taken to be an uncorrelated, uniformly distributed random variable, $\delta\mu(\mathbf{r})\in[-W,W]$. Shiba lattices could also present disorder in the adatom coupling strength, spin orientation, and possibly missing magnetic atoms in a given geometry [55]. In this study, we do not calculate the superconducting order parameter self-consistently in the presence of disorder [65]. In Fig. 7, we consider the disorder strength $W/t=1.2$ and compute the topological markers and local gaps for an island and a chain geometry. We average over a set of 100 disorder realizations. Panels (a) and (c) show the 2D and 1D local gaps, respectively. The dots show the average while the shaded area shows ±$\pm 1$ standard deviations. For this disorder strength, the 2D local gap is closed along the position $x$ path considered. The system is therefore 2D-trivial. However, the 1D local gap remains open, and system remains 1D-topological, as shown in panel (d), which shows $z_{2,\bm{\lambda}}$.
@@
Now, we consider the effect of disorder in the case where 1D and 2D topologies coexist in the clean case, as shown in Fig. 5, panels (n) and (s). We consider a disordered three-atom-tall Shiba island with disorder strength $W/t=0.3$ and average over 100 disorder realizations. The results are shown in Fig. 8. Panel (a) shows the 2D localizer gap. The solid gray line corresponds to the clean case, the blue dots to the average localizer gap, and the shaded area indicates ±$\pm 1$ standard deviations. The corresponding local Chern number, $|C_{\bm{\lambda}}^{\textrm{L}}|$, is shown in (b). The small gap protecting the topology in the clean limit is now significantly reduced. Panel (c) shows the 1D localizer gap, which remains open across the region with nontrivial $z_{2,\bm{\lambda}}=-1$. Therefore, the chemical potential disorder favors the 1D topology.
In the next section, we present the conclusions of our study and discuss possible extensions of our work to other physical systems and other symmetry classes.
@@
Figure 8: (a) 2D localizer gap for a disordered three-atom-tall Shiba island, as shown in the inset, with along position $x$ and $y=N/2$, with disorder strength $W/t=1.2$. The dots show the average over 100 disorder realizations, while the shaded area corresponds to ±$\pm 1$ standard deviations. Panel (b) shows the corresponding local maker, $|C_{\bm{\lambda}}^{\textrm{L}}|$, with orange squares. The clean limit results are shown in gray. (c) 1D localizer gap for a disordered Shiba chain along position $x$ and $y=N/2$, also for $W/t=1.2$, corresponding to the chain’s position. Panel (d) shows the corresponding local marker, $z_{2,\bm{\lambda}}$, which remains topological. The gray lines correspond to the clean limit. The spectral localizer calculations use $\kappa=0.05(t/l)$.
V Conclusions and outlook

### v2

@@ II.2 Spectral localizer
The 1D local topological invariant in class D is defined by first transforming to a basis for which $H$ is purely imaginary and $X$ is real. Such a basis must exist due to the system’s particle-hole symmetry. In this basis for $H$, and noting the choice of Pauli matrices used in Eq. (7), the 1D spectral localizer is both block off-diagonal and real symmetric at $E=0$, such that $(L^{(1D)}_{(x,0)})^{\dagger}=(L^{(1D)}_{(x,0)})^{\top}=L^{(1D)}_{(x,0)}$. Thus, all of its eigenvalues come in pairs, and when $L^{(1D)}_{(x,0)}$ is invertible, its matrix homotopy can be classified using only one of its off-diagonal blocks (which removes this spectral pairing) as [36]
$z_{2,x}=\text{sign}\left[\text{det}\left\{\kappa(X-xI)-iH\right\}\right].$ (8)
As the determinant is given by the product of a matrix’s eigenvalues, $z_{2,x}$ cannot change its value (for fixed $H$) without first crossing through a location in position-space where $\mu_{(x,0)}=0$. This reinforces the fact that $\mu_{\bm{\lambda}}$ provides both a measure of robustness as well as guarantee of bulk-boundary correspondence regardless of dimension and symmetry class. Note though, that locations where $\mu_{\bm{\lambda}}(X,H)=0$ are not generally connected to locations where $\mu_{\bm{\lambda}}(X,Y,H)=0$.
[…]
II.5 Shiba lattice dimensional crossover phase diagram
@@
Leveraging the local topological invariants for Class D systems, we explore the Shiba lattice’s parameter space for 1D and 2D geometries and consider values previously used in the literature for this model [72, 61, 73, 62]. To this end, we consider two geometries, a circular island, and a chain, and compute the corresponding local topological invariants (the Chern number $C_{\bm{\lambda}}^{\textrm{L}}$ and $z_{2,x}$, respectively) at the system’s center, see Fig. 3. Using the same parameters for both geometries, we find regions where the Shiba island and chain are both topological, regions where the Shiba island is topological and the chain is trivial, regions where the island is trivial and the chain is topological, and regions where both geometries are trivial. Note that topological transitions not captured in the diagram can occur as the geometry changes from 2D to 1D. For example, the transition from $|C_{\bm{\lambda}}^{\textrm{L}}|=2$ to $z_{2,x}=-1$ as the circular island is flattened occurs via a change to $|C_{\bm{\lambda}}^{\textrm{L}}|=1$ for a narrow island, as we discuss later [see Sec. II.6 and Fig. 6]. In Fig. 3, we consider only positive values for the chemical potential $\mu$, as the phase diagram is symmetric under $\mu\rightarrow-\mu$. In Supplementary Note 2, we present a lower-resolution phase diagram that considers negative values for the chemical potential $\mu$.
Even for the relatively modest system sizes considered in Fig. 3, the predicted topological phases of the Shiba island and chain are robust, as the localizer gaps [Figs. 3b,c,e,f] in both dimensions attain values within an order of magnitude of half the system’s bulk spectral gap [Figs. 1e,f]. Note, in an infinite clean system, the localizer gap at mid-gap is expected to be approximately half the bulk spectral gap width, i.e., the distance in energy to the nearest bulk states, as that is the perturbation to $L_{\bm{\lambda}}$ needed to find an approximate eigenstate. Indeed, there is a broad range of $\kappa$ spanning multiple orders of magnitude that produce quantitatively similar results, see Supplementary Note 3. Thus, the phase diagrams in Fig. 3a,d are also robust to the choice of position where the topology is determined, so long as this choice is a few sites from the Shiba geometry’s boundary.
II.6 Topological phase transitions from dimensional crossover
@@
Figure 7: Topological dimensional crossover as a function of island width. Length of the cut along $x$ for $y=N/2$ and $E=0$, as shown in Fig. 4, for which the local topological invariants $C^{\textrm{L}}_{\bm{\lambda}}$ (orange squares) and $z_{2,x}$ (black triangles) are nontrivial and localizer gap of appropriate dimension is larger than $0.02t$, $d^{\textrm{top}}$, divided by the number of sites along the cut, $d^{\textrm{tot}}$. a Lattice uses the parameters from Fig. 4. b Lattice uses the parameters $N=20$, $\mu=0$, $\alpha=0.8t/l$, $\Delta=1.2t$, and $JS=2t$. The set of plots showing the topological transition can be found in Supplementary Note 5. c Lattice uses the parameters from Fig. 6. The spectral localizer calculations use $\kappa=0.05(t/l)$.
To understand the various phases revealed in Fig. 3, we study the topological invariants as the Shiba geometry transitions from two-dimensional island to a one-dimensional chain by varying the distribution of magnetic adatoms placed atop the underlying superconductor.
@@
To begin, we consider the case where both the Shiba island and chain are topological with $|C_{\bm{\lambda}}^{\textrm{L}}|=1$ and $z_{2,x}=-1$, respectively, see Fig. 4. Specifically, our goal is to connect the change in the distribution of the magnetic adatoms [Fig. 4a-e] with the spectra of the associated spectral localizers [Fig. 4f-j], as well as the local markers and associated localizer gaps [Fig. 4k-t], as the adatom distribution is interpolated between a Shiba circular island and a Shiba chain. For the Shiba island, Fig. 4f, $\Sigma(L^{(2D)}_{\bm{\lambda}}(\bm{X},H))$ crosses zero at the edge of the magnetic adatom island, transitioning from trivial $C_{\bm{\lambda}}^{\textrm{L}}=0$ to topological with $|C_{\bm{\lambda}}^{\textrm{L}}|=1$ [Fig. 4k]. The finite localizer gap is maximal at the center of the island. However, for this same geometry, the spectrum of the 1D spectral localizer, $\Sigma(L^{(1D)}_{\bm{\lambda}}(\bm{X},H))$, displays a much smaller localizer gap within the Shiba island [Fig. 4p]. While there are sites where the topological local marker is topological, $z_{2,x}=-1$, the vanishing localizer gap means that there is little protection against variations in the Hamiltonian $\delta H$ or position operators $\delta X$. Therefore, the 1D topology is not robust, and the system is characterized by a two-dimensional topology via the local Chern number, $C_{\bm{\lambda}}^{\textrm{L}}$.
@@
As the number of magnetic adatoms is reduced along the y-direction, flattening the island, as shown in Fig. 4b, similar trends in the local markers and the associated localizer gaps are obtained. However, the number of sites with a nontrivial Chern number is reduced compared with the circular island’s case. Additionally, the 2D localizer gap is now smaller. The trend continues for a flatter island, as shown in Fig. 4c. However, for the parameters considered, the three-atom-wide island exhibits a change in its behavior: the Chern number is now nonzero only at locations where the 2D localizer gap vanishes, Fig. 4n, while there are sites with a nontrivial $z_{2,x}$ invariant protected by a 1D localizer gap, Fig. 4s. As such, this marks the transition from 2D to 1D topology in the system due to the change in the system’s geometry. For the Shiba chain, panel (e), the $z_{2,x}$ invariant changes from trivial to nontrivial at the edge of the chain. As the Shiba island reduces to a chain, $\Sigma(L^{(2D)}_{\bm{\lambda}=(x,y=N/2,E=0)}(\bm{X},H))$ approaches $\Sigma(L^{(1D)}_{\bm{\lambda}=(x,E=0)}(\bm{X},H))$. The local density of states for all the geometries considered is shown in the Supplementary Note 4.
To demonstrate that the topological transition shown in Fig. 4 is robust regardless for different boundaries of adatoms (so long as additional symmetries are not introduced), we consider the dimensional crossover using a geometry without symmetries, as shown in Fig. 5. To define the asymmetric Shiba islands, we start from the equation of a circle and define a “local” radius by introducing a random variable that changes the distance from the center of the island to the edge by up to $10\%$. The same procedure is applied to the flattened geometries. For the Shiba lattices that are 2D-like structures, Fig. 5a,b, the topology is characterized by the local Chern number, $|C^{L}_{\bm{\lambda}}|$ and protected by the large 2D localizer gap. Similar to Fig. 4, for these 2D geometries, the 1D localizer gap is too small to protect a meaningful $z_{2,x}$. Conversely, as the randomized-boundary system approaches the Shiba chain limit, Fig. 5c-e, it becomes trivial from the 2D topological invariant perspective, while displaying regions with a nontrivial $z_{2,x}$ protected by a 1D localizer gap.
@@
We now consider a set of parameters that leads to a topological island with $|C_{\bm{\lambda}}^{\textrm{L}}|=2$ transitioning into a topological chain $z_{2,x}=-1$, see Fig. 6. For this case, there are two main differences with respect to the previous example we discussed. First, there is a topological transition from $|C_{\bm{\lambda}}^{\textrm{L}}|=2$ to $|C_{\bm{\lambda}}^{\textrm{L}}|=1$ when the Shiba island shrinks from seven to five atoms wide in the $y$ direction. Second, for the three-atom-wide island, there is a coexistence of non-trivial 2D and 1D topology with $|C_{\bm{\lambda}}^{\textrm{L}}|=1$ and $z_{2,x}=-1$, respectively, though both topological phases are only weakly protected due to the relatively small localizer gaps in both cases.
@@
In Supplementary Note 5, we show plots for a topological island with $|C_{\bm{\lambda}}^{\textrm{L}}|=2$ transitioning into the trivial chain, $z_{2,x}=1$. In this case, the number of sites with nontrivial topology decreases together with the localizer gap protection, until the system becomes trivial in both 1D and 2D. Finally, in Supplementary Note 6, we consider the opposite case of a trivial Shiba island with $|C_{\bm{\lambda}}^{\textrm{L}}|=0$ transitioning into a topological chain with $z_{2,x}=-1$. In this case, the topological phase transition due to dimensional crossover occurs for the three-atoms-wide island.
To conclude the discussion of topological phase transitions due to dimensional crossover in the clean limit, we consider a different approach to understanding these transitions based on the number of sites in the system exhibiting a non-trivial local marker with at least a moderate localizer gap, see Fig. 7. In particular, we consider the number of sites with a topological local marker, $d^{\textrm{top}}$, divided by the total number of sites on the $x$-axis, $d^{\textrm{tot}}$, both for 1D and 2D, as a function of the island’s width-to-length ratio. Here, we only consider sites topological if the localizer gap is larger than $0.02t$. Thus, the transition from 2D topology to 1D topology, as characterized by the spectral localizer framework, is approximately given by the crossing of the traces.
@@
Figure 8: Effects of disorder on the Shiba island and chain. a Ensemble-averaged 2D localizer gap for a disordered circular Shiba island, as shown in the inset with filled and empty circles indicating magnetic adatom locations and absences, respectively, along $x$ for $y=N/2$ and $E=0$, and with disorder strength $W/t=1.2$. The blue dots show the average over 100 disorder realizations, while the light blue shaded area corresponds to $\pm 1$ standard deviation of the average. b Ensemble-averaged local maker $|C_{\bm{\lambda}}^{\textrm{L}}|$ (orange squares) and the region falling within $\pm 1$ standard deviation (light orange). c Ensemble-averaged 1D localizer gap (blue dots) and region falling within $\pm 1$ standard deviation (light blue) for a disordered Shiba chain, as shown in the inset, along $x$, for $E=0$ and $W/t=1.2$, corresponding to the chain’s position. d Ensemble-averaged local marker $z_{2,x}$ (orange squares) and the region falling within $\pm 1$ standard deviation (light orange). The spectral localizer calculations use $\kappa=0.05(t/l)$.
II.7 Effects of disorder
@@
To test the robustness of our results, we consider the effects of adding disorder in the local chemical potential. To do so, we include an on-site disorder term in the Hamiltonian,
$H_{\textrm{dis}}=-\sum_{\mathbf{r},\sigma}\delta\mu(\mathbf{r})c^{\dagger}_{\mathbf{r}\sigma}c_{\mathbf{r}\sigma},$ (9)
@@
where $\delta\mu(\mathbf{r})$ is taken to be an uncorrelated, uniformly distributed random variable, $\delta\mu(\mathbf{r})\in[-W,W]$. Shiba lattices could also present disorder in the adatom coupling strength, spin orientation, and possibly missing magnetic atoms in a given geometry [61]. In this study, we do not calculate the superconducting order parameter self-consistently in the presence of disorder [74]. Here, we consider the disorder strength $W/t=1.2$ and compute the topological markers and localizer gaps for an island and a chain geometry, averaging over an ensemble of 100 disorder realizations see Fig. 8. For this disorder strength, the 2D localizer gap is closed within the Shiba island along the path considered. Thus, whatever local 2D marker the system may possess, the associated phase is not robust. However, the 1D localizer gap remains open and system remains 1D-topological, see Fig. 8d.
@@
Now, we consider the effect of disorder for a narrow Shiba lattice where 1D and 2D topologies coexist in the clean case, as considered in Fig. 6n,s. In particular, we study a disordered three-atom-wide Shiba island with disorder strength $W/t=0.3$ and average over an ensemble of 100 disorder realizations, see Fig. 9. When considering the system’s 2D topology, the small gap protecting its topological phase in the clean limit is now significantly reduced in the presence of disorder. However, again, the 1D localizer gap remains open across the region with nontrivial $z_{2,x}=-1$. Together, these results indicate the chemical potential disorder favors the 1D topology.
@@
Figure 9: Effects of disorder at the topological dimensional crossover point. a Ensemble-averaged 2D localizer gap for a disordered three-atom-wide Shiba island, as shown in the inset with filled and empty circles indicating magnetic adatom locations and absences, respectively, along $x$ for $y=N/2$ and $E=0$, and with disorder strength $W/t=0.3$. The blue dots show the average over 100 disorder realizations, while the light blue shaded area corresponds to $\pm 1$ standard deviation of the average. b Ensemble-averaged local maker $|C_{\bm{\lambda}}^{\textrm{L}}|$ (orange squares) and the region falling within $\pm 1$ standard deviation (light orange). c Ensemble-averaged 1D localizer gap (blue dots) and region falling within $\pm 1$ standard deviation (light blue) for a disordered Shiba chain along $x$, for $E=0$ and $W/t=0.3$, corresponding to the chain’s position. d Ensemble-averaged local marker $z_{2,x}$ (orange squares) and the region falling within $\pm 1$ standard deviation (light orange). In a-d, the gray lines correspond to the clean limit. The spectral localizer calculations use $\kappa=0.05(t/l)$.
III Discussion